OpenStudy (anonymous):

The square of a number exceeds that number by 12. What are the two possible solutions? 3 or -4 3 or 4 -3 or 4

OpenStudy (anonymous):

i THINK its -3 or 4 :)

OpenStudy (anonymous):

Thanks do you know what the equation is to get the solution?

OpenStudy (anonymous):

u have to solve x+12=x^2 to get the answer :)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

ure welcome! :)

OpenStudy (radar):

Just to be sure you did not stray off track. Hally was correct in her equation\[x ^{2}=x+12\]\[x ^{2}-x-12=0\]Factors to (x-4)(x+3)=0 x=4 x=-3 check them 4 squared is 16 which is 12 more than 4 good answer. -3 squared is 9 which is 12 more than -3 good answer Good work Hally you deserve a medal.

OpenStudy (anonymous):

Thanks also